OpenStudy (anonymous):

Find the equation of the perpendicular bisector of the line segment with endpoints (-6,-6) and (3,6)

OpenStudy (campbell_st):

find the midpoint of the line segment x = (-6 + 3)/2 y = (-6 + 6)/2 the midpoint is (-1.5, 0) the bisector passes through the midpoint. next find the gradient of the line m = rise/run m = ( 6 - -6)/(3 - - 6) m = 12/9 m = 4/3 so the slope is of the line segment is 4/3 then the perpendicular line segment is - 3/4 the equation of the line is y = -3/4x + b substitute the values from the midpoint to evaluate b and get the equation of the line.